Tutor can you check my work?
Solve log 36 6=x (the 36 is smaller and below the log)
1:6/36=0.66666666667
log 36 6=1
thank you

Answer by stanbon(48535)

log 36 6=x
This says "x is the power of 36 that gives you 6".
36^x = 6
6^(2x) = 6
2x = 1
x = 1/2
===========
Cheers,
Stan H.